OpenStudy (anonymous):

An object has a constant acceleration of 40 ft/sec2, an initial velocity of −20 ft/sec, and an initial position of 10 ft. Find the position function, s(t), describing the motion of the object. Please help ill give medal

OpenStudy (nuttyliaczar):

Do you remember the general position function?

OpenStudy (anonymous):

d(t)= integral v(t)dt= double integral a(t)?

OpenStudy (nuttyliaczar):

That's true but it only helps if you know one of those equations. The position function is: \[x _{0}+(dx/dt)t+(1/2)(d ^{2}x/dt ^{2})t ^{2}\] Or as you probably like to remember it: \[x _{0}+vt+\frac{ at ^{2} }{ 2 }\]

OpenStudy (nuttyliaczar):

But do keep in mind what you said too. It saves you the trouble of blindly memorizing three equations, and so you only need one of them to get the other two.

OpenStudy (nuttyliaczar):

Do you know what to do now? Because from here you just need to plug it in

OpenStudy (anonymous):

Not sure what to plug in

OpenStudy (anonymous):

just confused on the steps on how to solve the problem

OpenStudy (nuttyliaczar):

x0 is the initial position. v is the initial velocity (in retrospect I forgot to write v0 to be consistent). a is the acceleration. And t is time.

OpenStudy (nuttyliaczar):

I think you just need to set up the equation and leave t as is

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

so what would s(t) be or look like?

OpenStudy (anonymous):

would it be 10+ (-20)=40(t)^2/2?